OpenStudy (anonymous):

Find the volume of the square pyramid shown. Round to the nearest tenth if necessary.

OpenStudy (anonymous):

timed exam please help

OpenStudy (anonymous):

meeeerp

OpenStudy (anonymous):

http://www.mathsteacher.com.au/year10/ch14_measurement/25_pyramid/Image5484.gif see this V=1/3Ah where A is are of square and h is height

OpenStudy (nali):

V= (b^2 * h) / 3 here is the equation just plug the base and height in and u will be done

OpenStudy (anonymous):

answer would be around 605 check it

OpenStudy (anonymous):

whats ^ mean ?

OpenStudy (nali):

to the power of Example b^2 means b to the power of 2

OpenStudy (anonymous):

so squared

OpenStudy (anonymous):

\[V=1/3(a ^{2}\times h)\] a is side of square and h is perpendicular height solve this to get answer V=1/3 11*11*15

OpenStudy (nali):

yes

OpenStudy (anonymous):

605 is the answer thanks

OpenStudy (anonymous):

it is